How do you write 10 to the 5th power in standard form?
solution of the equation 3y-5y+10=36?
Find p: 2p+10=40?

Answer by checkley71(8403)

10^5=100000
------------------------
3Y-5Y+10=36
-2Y=36-10
-2Y=26
Y=26/-2
Y=-13 ANSWER.
-----------------------------------------------
2P+10=40
2P=40-10
2P=30
P=30/2
P=15 ANSWER.